OpenStudy (anonymous):

$$ \text { If } \large f(x)= \frac { (x+ \frac 1x)^6 + (x^6+ \frac 1{x^6})-12}{(x+ \frac 1x)^4 + (x^3+ \frac 1{x^3})} $$ $$\text{ and x > 0, find the maximum value of f(x) }$$

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

Calculus free please!

OpenStudy (anonymous):

Else it would be too tedious..

OpenStudy (anonymous):

lol I won't even dare to do it with Calculus

OpenStudy (anonymous):

Me 2 :D

OpenStudy (anonymous):

just plug it in wolfram, and let it compute the maximum!

OpenStudy (anonymous):

3 is the answer

OpenStudy (anonymous):

My calculator is crunching for the solution.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Wow , you guys are really lethargic :P

OpenStudy (anonymous):

How am I even supposed to start this problem, Hit and Trial would be useless

OpenStudy (anonymous):

yes it will be I suppose some sort of trying to reduce as perfect squares then cancelling might help ..

OpenStudy (anonymous):

try using substitution

OpenStudy (anonymous):

Hey I think I got it please don't post the solution

OpenStudy (anonymous):

I think the whole problem is a red herring

OpenStudy (anonymous):

there is no global maximum... but how do we show that?

OpenStudy (amistre64):

this one?

OpenStudy (anonymous):

there are minima tho

OpenStudy (anonymous):

at x = 1, the expression is 3....

OpenStudy (anonymous):

now shower me with medalz!!!

OpenStudy (anonymous):

is 3 the right answer

OpenStudy (anonymous):

3 is the right answer.... for the global minimum when x > 0... and not for the maximum (which does not really exist)

OpenStudy (amistre64):

this resembles x^6/x^4: x^2 to me .... but thats only at large values of x which means there is no max since x^2 goes to infinity right?

OpenStudy (anonymous):

right @amistre

OpenStudy (anonymous):

but why won't you let us use calculus to figure that out :(

OpenStudy (amistre64):

and what defines calculus?

OpenStudy (anonymous):

check this out, at x =1 denominator attains minima so it is a possible answer

OpenStudy (anonymous):

calculus is the field of professional mathematics which involves coming up with new integration rules....

OpenStudy (anonymous):

@stom and the value of the function at x=1 is 3 :-D

OpenStudy (anonymous):

i know its 3 at x=1, thats why i said 3 is the answer,

OpenStudy (anonymous):

at the very bottom it shows the global minimum, and a little before that it becomes quite apparent that there is no maximum.

OpenStudy (anonymous):

i didn't get it :( is hit and trial the only way

OpenStudy (anonymous):

there is no maxima ,

OpenStudy (anonymous):

take t = x+1/x

OpenStudy (anonymous):

\[\frac{\left(x + \frac{1}{x}\right)^6 + \left(x + \frac{1}{x}\right)\left[\left( x + \frac{1}{x}\right)^2 -3\right] - 14}{\left( x + \frac{1}{x}\right)\left[\left(x + \frac{1}{x}\right)^3 + \left(x + \frac{1}{x}\right)^2 - 3\right]}\] This is as far as I can get :-/

OpenStudy (anonymous):

Whoopsy \[\frac{\left(x + \frac{1}{x}\right)^6 + \left(x + \frac{1}{x}\right)^2\left[\left( x + \frac{1}{x}\right)^2 -3\right]^2 - 14}{\left( x + \frac{1}{x}\right)\left[\left(x + \frac{1}{x}\right)^3 + \left(x + \frac{1}{x}\right)^2 - 3\right]}\]

OpenStudy (anonymous):

so what is the correct way to solve this @ffm?

OpenStudy (anonymous):

Yeah I am curious too